a beach ball holds 800 cubic inches of air. another beach ball has a radius that is half that of the larger ball. how much air does the smaller ball hold?

Found 3 solutions by stanbon, edjones, mananth:
Answer by stanbon(75887)

a beach ball holds 800 cubic inches of air. another beach ball has a radius that is half that of the larger ball. how much air does the smaller ball hold?
------
Radius of 800 cubic in ball = ?
(4/3)pir^3 = 800
r^3 = 190.99 in
r = 5.759 in
----
Vol. of smaller ball
V = (4/3)pi(5.759/2)^3
---
V = 100 cu inches
======================
Cheers,
Stan H.

Answer by edjones(8007)

V(sphere)=4/3 * pi * r^3
4/3 * pi * r^3=800
pi*r^3=600
r^3=600/pi
r=5.7588
5.7588/2=2.8794
When r=2.8794 then 4/3 * pi * r^3=100 in^3
.
Ed

Answer by mananth(16946)

the volumes of air in the two beach balls have a ratio of 8 times their volumes if the ratio of radii = 2
...
800/8
= 100 cu.inches
...
m.ananth@hotmail.ca